OpenStudy (anonymous):

Write the equation for the simple harmonic motion of a ball on a spring that starts at its lowest point of 6 inches below equilibrium, bounces to its maximum height of 6 inches above equilibrium, and returns to its lowest point in a total of 2 seconds.

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

\[ m {d^2 x \over dt^2} = - kx, x(0) = x(2) -6, \]

OpenStudy (experimentx):

the time period is 2 ... the solution should be harmonic function. so the answer should be \[ x(t) = -6 \cos \left( \pi t\right) \]